OpenStudy (dolphin1998):

Lara wrote the statements shown in the chart below: Statement Description 1 If two lines intersect, then they intersect at exactly one point. 2 In a right triangle, the square of the length of the hypotenuse is equal to the sum of the squares of the length of the legs. Which option best classifies Lara's statements? Select one: a. Statement 1 and Statement 2 are postulates because they are true facts. b. Statement 1 is a theorem, and Statement 2 is a postulate. c. Statement 1 is a postulate, and Statement 2 is a theorem. d. Statement 1 and Statement 2 are theorems because they

OpenStudy (anonymous):

I would say Statement 1 is a theorem because a line can lie on top of another line, where it has infinite solutions. I would say Statement 2 is a postulate because it is true. a^2+b^2=c^2.

OpenStudy (dolphin1998):

Okay how did you get that????

OpenStudy (just_one_last_goodbye):

Interesting isn't it?

OpenStudy (anonymous):

Well if you know what a theorem and a postulate is, you should understand it. A theorem is a rule in algebra or other branches of mathematics expressed by symbols or formula. A postulate is a statement that is accepted without proof. And a^2+b^2=c^2 is the formula for pythagorean theorem (which i got from the hypotenuse).

OpenStudy (dolphin1998):

yes